# Do We Need Large VLMs for Spotting Soccer Actions?

### v1

@@ Abstract
Traditional video-based tasks like soccer action spotting rely heavily on visual inputs, often requiring complex and computationally expensive models to process dense video data. In this work, we propose a shift from this video-centric approach to a text-based task, making it lightweight and scalable by utilizing Large Language Models (LLMs) instead of Vision-Language Models (VLMs). We posit that expert commentary, which provides rich, fine-grained descriptions and contextual cues such as excitement and tactical insights, contains enough information to reliably spot key actions in a match. To demonstrate this, we use the SoccerNet Echoes dataset, which provides timestamped commentary, and employ a system of three LLMs acting as judges specializing in outcome, excitement, and tactics. Each LLM evaluates sliding windows of commentary to identify actions like goals, cards, and substitutions, generating accurate timestamps for these events. Our experiments show that this language-centric approach performs effectively in detecting critical match events, providing a lightweight and training-free alternative to traditional video-based methods for action spotting.
@@ 1 Introduction
In the domain of video understanding, visual frames have traditionally been considered the best input for many tasks, including action spotting, event detection, and object recognition. However, these methods often require significant computational resources to process and analyze the dense video data. Despite the advancements in video models, such as convolutional neural networks (CNNs) and vision transformers (ViTs), the need for high-resolution video inputs can be prohibitive in both training and deployment scenarios. In many video contexts, particularly in sports, there is a spoken commentary that accompanies the visual content. This commentary provides detailed descriptions of the ongoing actions, events, and emotions, offering an alternative source of information. It is intriguing to consider whether these spoken descriptions, which can often provide rich contextual cues, could replace or complement visual inputs for specific tasks. By focusing on the textual information provided by commentators, we may gain a lightweight, scalable solution that avoids the complexity of video-based models.
Action spotting Seweryn et al. (2023), a core task in sports analytics, aims to identify key events within a video, such as goals, penalties, or substitutions, by analyzing the visual content. Traditional approaches have relied on object detection and tracking techniques that require parsing every frame of the video to detect specific actions. These methods can be computationally expensive and often struggle with long sequences or multiple simultaneous events. In contrast, when considering the commentary, each moment in the match is often described in rich detail, including the action, the players involved, and the contextual relevance. The spoken word can provide a nuanced understanding of the match dynamics, capturing moments of excitement, controversy, and strategic importance that may not always be fully conveyed through visual data alone. This raises an interesting possibility: Can we leverage textual commentary as a primary input for action spotting, bypassing the need for video frames?
In this paper, we explore this question by converting the traditional video-based action spotting task into a text-based one. Specifically, we investigate whether expert commentary, which provides fine-grained descriptions and contextual cues, is sufficient for reliably spotting key actions. The use of Large Language Models (LLMs) enables us to process and evaluate these textual descriptions, which contain detailed insights into match events, without requiring visual data. This approach not only alleviates the need for computationally intensive video models but also opens up new avenues for action detection in contexts where visual data may be scarce or of low quality, such as archived footage or radio broadcasts.
@@
We define two key research questions (RQs) for our study:
Research Question 1 Can fine-grained commentary provide sufficient information for accurate action spotting? In this study, we investigate whether the detailed descriptions provided by commentators are enough to reliably detect key actions, such as goals, cards, and substitutions, without relying on visual data.
Research Question 2 How can we leverage Large Language Models (LLMs) for event detection in textual commentary? We examine how a system of three LLMs, each specialized in outcome, excitement, and tactics, can be used to evaluate timestamped commentary windows and identify event-worthy segments.
Through our study, we provide the following contributions:
  • We utilize the SoccerNet Echoes Gautam et al. (2024) dataset, which contains timestamped commentary of soccer matches, to reformulate action spotting as a text-based task.
  • We design and implement a three-LLM system that judges the commentary based on outcome (goals, cards), excitement (crowd reactions, dramatic moments), and tactics (substitutions, formations).
  • We demonstrate that fine-grained commentary, in many cases, provides comparable or even superior information for event detection compared to visual cues.
@@ 2 Related Work
Fine-Grained Descriptions in Video-Based Tasks.
In video-based understanding tasks, traditional models have primarily relied on visual features extracted from video frames to detect and classify events. However, recent research has begun exploring the use of fine-grained descriptions—specifically, textual information derived from transcriptions or commentary—to enhance performance in tasks like action spotting and event detection. Xie et al. (2019) demonstrated that integrating visual information with text can improve performance in action recognition tasks, as descriptive cues often convey context that is missed by raw visual data. Fine-grained descriptions, which often come from sources such as narrators or commentators, offer a complementary modality that is particularly useful in contexts where visual cues are ambiguous or incomplete. In addition, Su et al. (2012) highlighted the utility of crowd-sourced commentary to aid in object detection tasks, which suggests that action spotting in dynamic environments, such as sports, could be enhanced by considering detailed narrative descriptions. These works point toward the potential of substituting or supplementing visual cues with textual descriptions, thereby providing more precise context for video interpretation tasks.
@@
Action spotting in soccer has long relied on visual inputs, particularly tracking players and ball movements. However, recent developments in leveraging commentary and other textual sources for action detection have gained attention. Giannakopoulos et al. (2016) proposed a method that uses timestamped commentary as input to detect key moments in soccer, such as goals or penalties, demonstrating that textual data can complement traditional visual cues. Another approach by Andrews et al. (2024) used a multi-modal network that combines both video frames and textual commentary to detect key events in football matches. These methods, however, often require substantial computational resources and may struggle with temporal consistency between visual events and their descriptions. While these advancements are promising, they highlight the gap in efficiently using fine-grained commentary without relying heavily on computationally expensive visual inputs or needing large amounts of data for training.
Despite these promising advancements, there remains a gap in fully utilizing fine-grained commentary for video understanding tasks like action spotting, especially in the context of soccer. Existing methods either rely on computationally expensive visual cues or fail to achieve consistent performance with textual input alone. In this paper, we address this gap by proposing a novel method that employs a three-LLM (Large Language Model) judge setup for action spotting based purely on textual commentary.
@@ 3 Methodology
SoccerNet-Echoes Dataset.
The SoccerNet-Echoes dataset is an essential component of our approach. This dataset provides detailed English commentary for soccer matches, which is transcribed from video using Whisper-based Automatic Speech Recognition (ASR). Whisper Radford et al. (2022) is known for its ability to transcribe multiple languages and its robustness to different accents, making it suitable for real-time, dense captioning in soccer broadcasts. The transcriptions are then formatted as timestamped text, which is far more concise and contextually richer than the video-based dense captions often provided by traditional action spotting models. The commentary text provides rich contextual information that can guide the identification of meaningful game actions, such as goals, fouls, and tactical shifts, and is the primary input to our semantic judges.
@@
We use Llama 3.1 8B Grattafiori et al. (2024) to create three specialized semantic judges, each focusing on different aspects of the game, highlighted in Figure 1.Each judge is steered by a dedicated system prompt plus 2–3 few-shot exemplars.
[…]
Goals, penalties, red/yellow cards.
[…]
Crowd eruptions, near-misses, controversies.
[…]
Substitutions, formation changes, momentum swings.
Our classifier goes from binary to class-level matching, providing a quick look (Yes/No) for majority voting (next step), and a detailed classwise matching as highlighted below
Input: full English commentary for a 10 s window. Output: 1. YES/NO – does the window contain an action-worthy event? 2. If YES: classify into one of the 17 SoccerNet-V2 classes (Fig. 2) or label as “out-of-world” when relevant but off-taxonomy.
Intuition For many soccer matches, emotional moments such as goals and fouls are often accompanied by an excited tone in the commentary. This emotional context is a strong signal for the excitement judge, which can detect such moments without needing to rely on visual cues. The strategic context provided by the tactical judge is similarly informed by phrases that suggest changes in team strategy or formation.
Figure 2: A detailed look into the 17 action classes for Soccernet-v2.
@@ 4 Evaluation and Results
We evaluate the performance of our proposed method using two metrics: average mean Average Precision (mAP) and tight mAP. The results on Soccernet-v2 Deliege et al. (2021) are summarized in Table 1, show that our LLM-based method achieves a significant improvement over existing baselines. Specifically, our method achieves an average mAP of 64.50% and a tight average mAP of 60.75%, outperforming the SoccerNet baseline Cioppa et al. (2020) by a margin of 14.80% in average mAP and 15.75% in tight mAP. It also surpasses RMS-Net Tomei et al. (2021) by 11.10% in average mAP and 11.25% in tight mAP.
Commentary-Anchored Random Baseline.
@@
Results.
When compared to the state-of-the-art method, ASTRA Xarles et al. (2023), our method achieves 96.51% of the average mAP and 97.23% of the tight average mAP, bringing us very close to the top-performing model. This demonstrates that our approach is highly competitive, offering strong performance with a focus on leveraging text-based information. Our method was implemented on an RTX 3060 GPU.
Method mAP (%) Tight
mAP (%)
CALF Cioppa et al. (2020) 49.7
RMS-Net Tomei et al. (2021) 53.4
ASTRA Xarles et al. (2023) 79.2 70.1
Random Text-Only 12.0 10.5
LLM-Based (Ours) 64.5 60.8
Table 1: Comparison of our method with existing approaches. Note that both metrics refer to average mAP as discussed in the beginning of Section 4. Our text-only method comes close to parameter-heavy highly trained action spotters.
@@ Discussion.
Our current work raises the question if video frames, which are resource demanding in processing, needs to be included for action spotting. One can posit that this extends to a higher, more inclusive question: Do we need VLMs to spot any action where it is already being described in expert depth? Our current scope and focus in the paper lies in showing its capabilities for soccer match videos, which de-facto provide for expert-level realtime descriptions of events ongoing in the match. We notice compared to a blind lower bound baseline and the current SOTA VLM for action spotting, harnessing the abilities of few-shot LLMs is enough given the quality of commentary is rich. We plan to extend this to the aforementioned greater RQ and also provide a detailed study as to how quality of commentary affects any activity-based video action spotting based on our framework.

### v2

@@ Abstract
Traditional video-based tasks like soccer action spotting rely heavily on visual inputs, often requiring complex and computationally expensive models to process dense video data. We propose a shift from this video-centric approach to a text-based task, making it lightweight and scalable by utilizing Large Language Models (LLMs) instead of Vision-Language Models (VLMs). We posit that expert commentary, which provides rich descriptions and contextual cues contains sufficient information to reliably spot key actions in a match. To demonstrate this, we employ a system of three LLMs acting as judges specializing in outcome, excitement, and tactics for spotting actions in soccer matches. Our experiments show that this language-centric approach performs effectively in detecting critical match events coming close to state-of-the-art video-based spotters while using zero video processing compute and similar amount of time to process the entire match.
@@ 1 Introduction
In the domain of video understanding Nguyen et al. (2024), visual frames have traditionally been considered the best input for many tasks, including action spotting, event detection, and object recognition Giancola et al. (2025, 2023); Fulari (2018) . However, these methods often require significant computational resources to process and analyze the dense video data Selva et al. (2023); Feichtenhofer et al. (2019). Despite the advancements in video models, such as convolutional neural networks (CNNs) Karpathy et al. (2014) and vision transformers (ViTs), the need for high-resolution video inputs can be prohibitive in both training and deployment scenarios.
Action spotting Seweryn et al. (2023), a core task in sports analytics, aims to identify key events within a video, such as goals, penalties, or substitutions, by analyzing the visual content. Manual methods by broadcasters were slow and took time in distribution Merler et al. (2019).Traditional approaches Shih (2017) have relied on object detection and tracking techniques that require parsing every frame of the video to detect specific actions Khan et al. (2018). These methods can be computationally expensive and often struggle with long sequences or multiple simultaneous events Xu et al. (2025). In contrast, when considering the commentary, each moment in the match is often described in rich detail, including the action, the players involved, and the contextual relevance. The spoken word can provide a nuanced understanding of the match dynamics, capturing moments of excitement, controversy, and strategic importance that may not always be fully conveyed through visual data alone. This raises an interesting possibility: Can we leverage textual commentary as a primary input for action spotting, bypassing the need for video frames?
We explore this question by proposing a text based action spotting pipeline using an LLM-as-a-judge setup, following Zheng et al. (2023). We investigate whether expert commentary is enough for current LLMs to infer actions from, and if it is comparable to heavy video based action spotter VLMs. We also study the improvement in action spotting as time taken per match and the independence from video processing compute.To this end, we provide the following contributions:
@@
  • We redesign action spotting as a text based task as compared to a visual based task, utilising the Soccernet-Echoes dataset Gautam et al. (2024).
  • We design and implement a three-LLM system that judges the commentary based on outcome , excitement, and tactics.
  • We demonstrate that expert commentary, in many cases, provides comparable information for event detection compared to visual cues.
@@ 2 Related Work
Detailed Descriptions in Video-Based Tasks.
In video-based understanding tasks, traditional models have primarily relied on visual features extracted from video frames to detect and classify events. However, recent research has begun exploring the use of fine-grained descriptions, specifically, textual information derived from transcriptions or commentary, to enhance performance in tasks like action spotting and event detection. Xie et al. (2019) demonstrated that integrating visual information with text can improve performance in action recognition tasks, as descriptive cues often convey context that is missed by raw visual data.In addition, Su et al. (2012) highlighted the utility of crowd-sourced commentary to aid in object detection tasks, which suggests that action spotting in dynamic environments, such as sports, could be enhanced by considering detailed narrative descriptions. Recent work shows that textual descriptions can carry action semantics the pixels miss and when transcribed reliably, can act as a compact surrogate for frames. For soccer specifically, dense, timestamped commentary corpora like SoccerNet-Caption Mkhallati et al. (2023) and GOAL Qi et al. (2023) establish the feasibility of commentary-anchored modeling, while MatchTime Rao et al. (2024) highlights and fixes video-text misalignment—a key pain point for using commentary in downstream tasks. Robust ASR models such as Whisper Radford et al. (2022) makes multi-accent, broadcast-noise transcripts viable at scale, strengthening the case for text-first pipelines.
@@
Action spotting in soccer has long relied on visual inputs, particularly tracking players and ball movements. However, recent developments in leveraging commentary and other textual sources for action detection have gained attention. Giannakopoulos et al. (2016) proposed a method that uses timestamped commentary as input to detect key moments in soccer, such as goals or penalties, demonstrating that textual data can complement traditional visual cues. Another approach by Andrews et al. (2024) used a multi-modal network that combines both video frames and textual commentary to detect key events in football matches. The SoccerNet benchmark Deliege et al. (2021) formalized spotting as timestamp localization, driving a largely video-first literature . Classical baselines learn visual features and pool them temporally such as CALF Cioppa et al. (2020) and NetVLAD++ temporal pooling Giancola and Ghanem (2021) . Subsequent models improved localization via stronger heads/sequence learning, including RMS-Net Tomei et al. (2021) and compact E2E-Spot Hong et al. (2022). Recent transformer systems such as ASTRA Xarles et al. (2023) push tight-tolerance accuracy further and even add audio for non-visible cues. Broader universal efforts such as UniSoccer Rao et al. (2025) argue for richer taxonomies and multi-task foundations that still place video at the center. These threads collectively set a strong video baseline for action spotting.
Despite these promising advancements, there remains a gap in fully utilizing fine-grained commentary for video understanding tasks like action spotting, especially in the context of soccer. Existing methods either rely on computationally expensive visual cues or fail to achieve consistent performance with textual input alone.
@@ 3 Methodology
We use Llama 3.1 8B Grattafiori et al. (2024) to instantiate three specialized judges that operate over a shared label space of the 17 SoccerNet-V2 classes and NO-ACTION. Each judge sees the same 10 s commentary window (5 s stride) but is prompted with a distinct evidence lens (Outcome, Tactics, Emotion). All three judges return a single class (or NO-ACTION) and confidence score. Judges are steered by a dedicated system prompt and 2–3 few-shot exemplars .
Outcome-centric Judge
Prioritizes refereeable outcomes (goal, penalty, yellow/red), explicit referee phrases.
Tactics-centric Judge
Emphasizes set-pieces and structure (corner, free-kick, substitution, formation/press).
W Emotion-centric Judge
Uses rhetorical intensity and urgency to resolve ambiguous cases; conservative when negations appear (“over the bar”, “flag is up”).
Input: full English commentary for a 10 s window (5 s stride). Output (per judge):
1. A single label in {17 SoccerNet-V2 classes} $\cup$ {NO-ACTION}. 2. A confidence in $[0,1]$ (calibrated from model’s self-score).
Abstention is expressed as NO-ACTION; we use higher thresholds for the Emotion judge to avoid rhetorical over-triggering.
@@
Method M mAP (%) Tight mAP (%)
CALF Cioppa et al. (2020) Video 49.7
RMS-Net Tomei et al. (2021) Video 63.49 28.83
FCMA Zhou et al. (2021) Video 73.77 47.05
E2E-Spot (RegNetY-200MF) Hong et al. (2022) Video 73.25 61.19
E2E-Spot (RegNetY-800MF) Hong et al. (2022) Video 74.05 61.82
ASTRA Xarles et al. (2023) Video 78.09 66.82
Random Text-Only (ours, baseline) Text 12.0 10.5
LLM-Based (Ours) Text 64.5 60.8
Table 1: mAP and tight mAP on SoccerNet-v2 for video- vs text-based pipelines. M = Modality (Video/Text).
Method Input FLOPs/frame (GF) Total FLOPs (TF) Time/frame (ms) Time/match (sec)
RegNetY-200MF (E2E-Spot) Video 0.20 2.16 0.3 3.24
RegNetY-800MF (E2E-Spot) Video 0.80 8.64 0.9 9.72
ResNet-152 (baseline feats) Video 11.5 124.2 1.8 19.44
R(2+1)D (3D CNN) Video 11.0 118.8
Ours: LLM (text-only) Text 146.5
Table 2: Efficiency comparison on FLOPs and wall-clock GPU/CPU time needed for a full match evaluation.
@@ 4 Evaluation and Results
Setup.
We evaluate on the SoccerNet-v2 test split using the SoccerNet-Echoes commentary Gautam et al. (2024) as input. Events follow the 17-class SoccerNet taxonomy. Our system operates on 10 s windows (5 s stride) and uses three Llama 3.1 8B judges. We report the official average-mAP (“mAP”) over $\delta\!\in\![5,60]$ s and tight Average-mAP (“Tight”) over $\delta\!\in\![1,5]$ s in 1. For efficiency (Table 2) we normalize video compute to a 90 min match at 2 FPS (10,800 frames) and report backbone FLOPs/frame and measured per-frame time on an A5000 from prior work Hong et al. (2022) and for our text-only system we report CPU wall-clock time per match.
Textual Random Baseline.
@@
Main results.
[…]
Efficiency ablation.
Table 2 compares per–90 min match compute and explains our savings. Video pipelines pay a cost that scales with the number of visual tokens; text scales with text tokens. Let $F$ be frames per match and $P$ the patch tokens per frame (ViT-style). Then $\text{visual-tokens}=F\!\times\!P$ and $\text{text-tokens}=N_{t}$. At 2 FPS, F=10,800F{=}10{,}800. For ViT-B/16 at $224^{2}$, $P{=}(224/16)^{2}{=}196$, so $F\!\times\!P\!\approx\!2.12\times 10^{6}$ visual tokens/match, whereas ASR produces only $N_{t}=\mathcal{O}(10^{4})$ text tokens—two orders of magnitude fewer. Even with CNNs (no explicit patches), the effective per-frame compute (GFLOPs/frame) still scales with $F$ and dominates.
At 2 FPS, published video backbones span 2.16–124.2 TFLOPs per match and 0.3–1.8 ms per frame on an A5000 (3.24–19.44 s per match; a 3D CNN is 118.8 s). Our pipeline performs no video feature extraction (zero video FLOPs) and instead scales with $N_{t}$ and LLM tokens/s. On CPU, our measured end-to-end time for a full match is 146.5 s (2.44 min), removing the dominant frame-processing term and any GPU requirement.
@@ Discussion.
(1) Tight localization from text. When outcomes are spoken (“penalty given”, “booked”, “and it’s in”), the language signal is temporally sharp, explaining our proximity to video SOTA on Tight mAP. (2) Loose-gap sources. Non-verbal micro-events and terse restarts are under-described in commentary, which hurts loose recall and favors video. (3) Design effects. Confidence thresholds and majority voting suppress rhetorical false positives (near-misses), improving precision; temporal NMS converts overlapping window votes into a single timestamp per event. (4) Compute and deployment. Zero-frame processing plus competitive Tight mAP make the approach attractive for CPU-scale batch processing (clubs/broadcasters) and for low-cost inference at volume.